# Discretion in the Loop: Human Expertise in Algorithm-Assisted College Advising

### v1

@@ Abstract
In higher education, many institutions use algorithmic alerts to flag at-risk students and deliver advising at scale. While much research has focused on evaluating algorithmic predictions, relatively little is known about how discretionary interventions by human experts shape outcomes in algorithm-assisted settings. We study this question using rich quantitative and qualitative data from a randomized controlled trial of an algorithm-assisted advising program at Georgia State University. Taking a mixed-methods approach, we examine whether and how advisors use context unavailable to an algorithm to guide interventions and influence student success. We develop a causal graphical framework for human expertise in the interventional setting, extending prior work on discretion in purely predictive settings. We then test a necessary condition for discretionary expertise using structured advisor logs and student outcomes data, identifying several interventions that meet the criterion for statistical significance. Accordingly, we estimate that 2 out of 3 interventions taken by advisors in the treatment arm were plausibly “expertly targeted” to students using non-algorithmic context. Systematic qualitative analysis of advisor notes corroborates these findings, showing that advisors incorporate diverse forms of contextual information—such as personal circumstances, financial issues, and student engagement—into their decisions. Finally, we explore the broader implications of human discretion for long-term outcomes and equity, using heterogeneous treatment effect estimation. Our results offer theoretical and practical insight into the real-world effectiveness of algorithm-supported college advising, and underscore the importance of accounting for human expertise in the design, evaluation, and implementation of algorithmic decision systems.
1 Introduction
@@
As algorithm–assisted decision making becomes increasingly common in high-stakes public domains, there is growing recognition that evaluating these systems requires more than assessing model performance in isolation. In settings where algorithms inform—but do not replace—human decision-makers, understanding how human experts actually use algorithmic tools is essential to assessing their real-world impact and equity implications [29, 36, 59, 18]. For example, in college advising, an algorithm may identify students at risk of dropping out, but it is up to a human counselor to have a conversation with the student and suggest a course of action based on the student’s specific circumstances.
@@
Evaluating human decision making when assisted by algorithms is essential for the design, evaluation, and implementation of automated decision systems. When algorithms are embedded in human decision processes, it becomes critical to ask: How much of the decision-making success is attributable to the algorithm versus the human expert? What discretionary work do humans perform that the algorithm cannot? What additional, unobserved information are important for improving outcomes? These questions help reveal whether key contextual or non-algorithmic information needs to be explicitly collected by humans, or whether algorithmic systems should be redesigned to incorporate this information.
Understanding the balance between algorithmic and human contributions becomes crucial for expanding and generalizing the educational intervention to new contexts (e.g., a new insitution). While algorithms can be deployed at scale without capacity constraints, the human component may introduce scaling challenges (e.g., social science interventions that perform well in initial studies but do not perform well at scale) due to changes in the implementation, resource constraints[13], or staff training [37, 46]. If human discretion plays a significant role in improving outcomes, then expanding the intervention requires either increasing the capacity of the existing experts or recruiting similarly qualified personnel.
We explore these questions in the context of college advising, a setting in which human decisions are both consequential and resource-constrained. State universities face growing pressure to improve student success and close equity gaps, and algorithm–assisted advising has been proposed as one strategy to help institutions scale proactive support [62, 24]. Yet the real-world impact of such systems depends crucially on how advisors use them.
There has long been interest in how predictive algorithms influence student outcomes, with early work focusing on identifying at-risk students through early warning systems at secondary and tertiary levels [43, 10, 66, 22, 51]. More recent work questioned the actionability and real-world impact of these systems [47, 48], with observational studies finding weak or null effects [21, 56]. Empirical studies of the impact of predictive systems are often limited by challenges in data collection and study implementation. Analyzing the causal effect of an algorithm–assisted system requires (1) longitudinal data on predictions, interventions, and downstream outcomes, which are not often documented at the same time, and (2) randomization of treatment assignment, which can be difficult to implement in educational institutions.
@@
Additionally, quantitative data alone often fail to capture how practitioners interpret and act on algorithmic outputs. Quantification can abstract away the social context and judgment that shape real-world decisions [71]. Qualitative data—such as advisors’ meeting notes or child welfare caseworker casenotes—can provide insight into “street-level” discretionary work [69, 70] that is not reflected in structured data. Such data are rarely collected in tandem with quantitative measures, especially in education.
Our study leverages a unique opportunity to fill these gaps: we combine rich quantitative and qualitative data from a randomized controlled trial to investigate the role of human discretion in algorithm–assisted college advising. Specifically, we draw on data from the landmark MAAPS experiment at Georgia State University [2, 64, 65]—a multi-year randomized controlled trial of a data-driven educational intervention described as “intensive, proactive, technology-enhanced advisement and degree planning.” The goal of the MAAPS study was to identify at-risk students and administer timely interventions to support these students through both regular advising meetings and algorithmic early warning systems.
The MAAPS data analyzed in this paper includes comprehensive quantitative records of timestamped advisor interactions and student outcomes per semester. Additionally, advisors in the MAAPS program systematically logged interventions performed during their meetings with students (e.g., recommending courses or degree planning) and recorded qualitative comments on their conversation with the student and specific guidance. These qualitative comments, in conjunction with the quantitative outcomes, provide a unique window into human decision making alongside algorithmic tools.
@@
Using a novel mixed-methods approach combining theoretical modeling, statistical testing, and qualitative analysis, we investigate how advisors leverage non-algorithmic information in student interventions in the MAAPS experiment, and explore the broader implications of discretionary decision making in algorithm-assisted advising. In meetings with students, advisors learn information about students’ interests, struggles, and personalities, creating an information asymmetry [60] between human decision makers and algorithmic tools that are entirely reliant on administrative data. Advisors may use this non-algorithmically available context to expertly target their interventions to students, improving performance over any possible algorithm. Our central research question is: What is the evidence for advisor discretion in algorithm–assisted advising, and what is its role in improving student outcomes?
Our contributions are as follows:
@@
1. Theoretical framework for discretionary expertise in data-driven interventions. We develop a causal graphical framework to define and audit for human expertise in the interventional setting. Extending prior work on expertise in pure prediction tasks [5, 6], we define expertly targeted action in terms of three causal conditions (effectiveness, targeting, and heterogeneoty) and derive a statistical condition that is necessary but insufficient for proving expertise in interventions. We highlight a fundamental impossibility result in the interventional setting: discretionary expertise cannot be identified from observational data alone.
@@
2. Empirical evidence of expert discretion in data-driven advising. Using quantitative data from the MAAPS randomized trial, we test for discretionary expertise by examining whether specific advisor interventions predict student outcomes (persistence, GPA) conditional on algorithmically available features. We identify five such ‘expert’ interventions ($p<0.05$) through a conditional independence test of the statistical condition identified in our theoretical framework. Subsequent qualitative analysis of advisor comments show that advisors indeed incorporate non-algorithmic context, such as the student’s personal circumstances, or financial troubles, into the ‘expert’ interventions.
3. Impact of advising on student outcomes and equity. We explore differences in advising styles across advisors, finding qualitative evidence of variation (i.e., academic vs. holistic focus) but no statistically significant advisor-specific effects on outcomes. A power analysis suggests our study is underpowered to detect many forms of heterogeneity, underscoring the methodological challenges of equity evaluation in real-world deployments.
@@
These findings illustrate the critical role of human discretion in algorithm-assisted interventions and highlight the importance of human factors to the success of data-driven college advising.
@@ 2 Related Work
Causal inference techniques are critical for evaluating algorithm-assisted decisions and their downstream effects. Notably, Imai et al. [36] conducted the first randomized field experiment to evaluate how an algorithmic decision aid in pre-trial risk assessment affects human judgment and decision quality in the U.S. criminal justice system. Follow-up work by Ben-Michael et al. [12] focused on whether algorithm-assisted human decisions improved decision-making “accuracy” over counterfactual baselines. Our work shifts the focus from decision quality per se to the mechanisms of human discretion: we investigate whether, how, and when human decision-makers contribute additional, outcome-relevant expertise. This is especially relevant in the educational setting where advisors choose among multiple possible interventions that may improve student outcomes.
In the educational domain, while the importance of understanding downstream effects of algorithmic student risk assessment has long been recognized [10], there has been little randomized evaluation of algorithmic decision making, aside from the current work, to the best of our knowledge. Perdomo et al. [55] use a regression discontinuity design to estimate the causal effect of early warning system risk scores on student success in Wisconsin public schools, finding small effects but have limited insight into underlying mechanisms due to missing data on subsequent interventions. Liu et al. [47] introduce a causal framework showing that accurate outcome prediction is insufficient for selecting effective interventions to improve individual outcomes in education.
Green and Chen [29] examine how algorithmic risk scores influence human decision making in a lab study, finding that human evaluators exhibit selective adherence to algorithmic recommendations, often amplifying existing biases. Raji and Liu [59] discuss methodological challenges of evaluating algorithmic-assisted decisions, emphasizing the need for experimental designs that disentangle the potential cognitive biases of a human judge.
@@
The role of evaluator expertise and bias in algorithm-assisted decisions has been a central concern in multiple domains. Holstein and Doroudi [33] and Baker and Hawn [8] explore how biases in algorithmic predictions can propagate through human decision making, particularly in educational contexts. Feathers [23] also examines how race and other demographic factors influence algorithmic predictions in higher education, raising equity and fairness concerns. These findings align with broader debates on the unintended consequences of algorithmic interventions in advising and student support.
Li [45] further investigates how expertise in peer review influences evaluators’ reliance on algorithmic recommendations, showing that domain experts may either mitigate or exacerbate algorithmic biases depending on their pre-existing heuristics and decision-making frameworks. Likewise, Hoffman et al. [32] analyze how hiring managers integrate algorithmic predictions into their decisions, demonstrating that discretion can either counteract or reinforce algorithmic biases. Gillis et al. [28] show that whether the algorithm amplifies or reduces the bias of the human decision-maker, with respect to a protected class, depends on whether the algorithm uses the protected class as a feature.
@@
In addition to Alur et al. [5], a growing body of work studies how humans combine their own judgment with the algorithmic predictions and what scenarios lead to humans relying more on algorithms. Balakrishnan et al. [9] show that people often combine their judgments with algorithmic recommendations, increasing prediction errors by over-relying on algorithms when their own insights are more valuable and vice versa. Snyder et al. [73] find that humans rely on algorithms when the algorithms perform better than humans (on average) and when the humans are required to make decisions quickly.
@@ 3.1.2 Data
For all 1037 students (520 in treatment and 517 in control) who participated in the MAAPS study, the dataset includes their baseline characteristics at the time of enrollment to GSU (such as their demographics, high school GPA and their scores on college entry tests), as well as timeseries data on their educational outcomes updated each term (such as their GPA and credit accumulation), including their graduation outcomes. This data allows us to understand how discretionary decision making connects to students’ educational outcomes and to look for heterogeneity of treatment effects in their baseline features. A full list of dataset variables are in Appendix A.1.
@@ 3.2 Our Mixed-Methods Approach
Our mixed-methods analytical framework is designed to investigate the role of advisor discretion and expertise within the treatment group. Our motivation for integrating quantitative and qualitative methods is three-fold: (1) to develop a comprehensive understanding of advisor expertise in interventional decision making, where qualitative insights help interpret and contextualize quantitative findings; (2) to support an exploratory analysis of heterogeneous treatment effects, especially in cases where quantitative signals may be subtle or underpowered; and (3) to produce novel qualitative insights that can directly inform professional development for advisors, reveal gaps, and support policy recommendations.
@@
We begin by reflecting on our positionality as researchers in this study. We recognize that our identities, disciplinary training, and institutional affiliations shape how we approached the research questions, interpreted the data, and interacted with the broader context of advising and algorithmic decision making. The research team includes graduate students (Authors 1 and 2) and tenure-track faculty (Authors 5 and 6) at R1 institutions, with training in computer science, operations research, and the societal impacts of algorithmic decision making, including secondary expertise in educational applications.
The research process was shaped by an ongoing, collaborative partnership between the research team and staff at the National Institute for Student Success (NISS). From November 2023 to November 2024, Authors 5 and 6 held biweekly meetings with three partnering staff members at NISS, including Authors 3 and 4, to understand the MAAPS intervention and its implementation context, as well as to determine the structure and limitations of the advising data. These meetings informed both our quantitative and qualitative analyses by grounding them in the realities of the program.
Our access to the data was limited to de-identified records shared under a data use agreement with NISS, and we did not interact directly with students or advisors. The Princeton University and Columbia University Institutional Review Boards determined the research was exempt from further review, as it does not qualify as human subjects research. We recognize that our interpretation of advising practices is shaped by our outsider perspective and by the structure and limitations of the available data. In the qualitative analysis, we sought to preserve the nuance of advisors’ written reflections while acknowledging the constraints of interpreting them at a distance.
Finally, as this study performs a secondary analysis of an experiment designed primarily to estimate average treatment effects, the original experimental design and data availability have shaped both the scope of our research questions and methodological choices.
@@ 4 Theoretical Framework: Expertly Targeted Interventions
A student meets with their advisor, explaining that a medical condition prevents them from walking long distances. In response, the advisor works with the student to design a schedule with nearby classrooms that also meets her academic interests. The student is able to achieve a high semester GPA.

### v2

@@ Abstract
In higher education, many institutions use algorithmic alerts to flag at-risk students and deliver advising at scale. While much research has focused on evaluating algorithmic predictions, relatively little is known about how discretionary interventions by human experts shape outcomes in algorithm-assisted settings. We study this question using rich quantitative and qualitative data from a randomized controlled trial of an algorithm-assisted advising program at Georgia State University. Taking a mixed-methods approach, we examine whether and how advisors use context unavailable to an algorithm to guide interventions and influence student success. We develop a causal graphical framework for human expertise in the interventional setting, extending prior work on discretion in purely predictive settings. We then test a necessary condition for discretionary expertise using structured advisor logs and student outcomes data, identifying several interventions that meet the criterion for statistical significance. Accordingly, we estimate that 2 out of 3 interventions taken by advisors in the treatment arm were plausibly “expertly targeted” to students using non-algorithmic context. Systematic qualitative analysis of advisor notes corroborates these findings, showing a pattern of advisors incorporating diverse forms of contextual information—such as personal circumstances, financial issues, and student engagement—into their decisions. Our results offer theoretical and practical insight into the real-world effectiveness of algorithm-supported college advising, and underscore the importance of accounting for human expertise in the design, evaluation, and implementation of algorithmic decision systems.
@@
As algorithm–assisted decision making becomes increasingly common in high-stakes public domains, there is growing recognition that evaluating these systems requires more than assessing model performance in isolation. In settings where algorithms inform—but do not replace—human decision-makers, understanding how human experts actually use algorithmic tools is essential to assessing their real-world impact and equity implications [26, 33, 55, 16, 25, 19]. For example, in college advising, an algorithm may identify students at risk of dropping out, but it is up to a human counselor to have a conversation with the student and suggest a course of action that is tailored to the student’s specific circumstances. Despite the widespread deployment of such systems, empirical research remains limited on how human decision-makers interpret algorithmic recommendations, exercise discretion, and influence downstream outcomes.
The Monitoring Advising Analytics to Promote Success (MAAPS) experiment at Georgia State University provides a rare opportunity to investigate these questions in a real-world setting. This multi-year (2016-2020), randomized controlled trial evaluated an algorithm-assisted advising program that offered “intensive, proactive, technology-enhanced advisement and degree planning,” identifying students to support through both advising meetings and algorithmic early warning systems. Notably, the MAAPS program was evaluated via a carefully designed and well-documented experiment that demonstrated positive treatment effects on graduation and persistence [59, 60].
@@
While studies like the MAAPS experiment offer encouraging evidence that algorithmic early warning systems can improve student outcomes, they leave open important questions about why these interventions work. A recent review, for instance, notes that existing studies offer “largely speculative interpretations of intervention mechanisms and effects” [40]. One key question is about the role of human discretion. When algorithms are embedded in human decision processes, it becomes critical to ask: How much of the decision-making success is attributable to the algorithm versus the human expert? What discretionary work do humans perform that the algorithm cannot? What additional information do they need to affect outcomes? These questions help reveal whether non-algorithmic information needs to be explicitly collected by humans, or whether algorithmic systems should be redesigned to better support interventions for student success.
Recognizing that human discretion could shape the impact of algorithm-assisted decision making, our central research question is: What is the evidence for advisor discretion in algorithm-assisted advising, and what is its role in improving student outcomes? Understanding the balance between algorithmic and human contributions is key to scaling algorithm-assisted advising, a strategy that is increasingly used by public universities to boost student success and close equity gaps [57, 21]. While algorithms can be scaled without capacity constraints, the human component cannot — reflecting a hard-won lesson of the social sciences, that interventions performing well in initial studies may not scale due to changes in the implementation, available resources, or personnel training [34, 43]. If human discretion is a key driver of impact, scaling such systems will require not only algorithmic technologies but also investment in human expertise and skills.
To examine the role of human discretion in the success of algorithm-assisted advising, we analyze detailed data on advisor decision-making within the treatment group of the MAAPS trial. The dataset includes structured, time-stamped records of advisor interactions and student outcomes for each semester. In addition to these quantitative logs, MAAPS advisors systematically documented the interventions they delivered during meetings with students (e.g., recommending courses or degree planning) and recorded qualitative comments on their conversations.
@@
This mixture of qualitative and quantitative data provides a unique window into human decision making alongside algorithmic tools. Quantitative data alone often fail to capture how practitioners interpret and act on algorithmic outputs, and can abstract away the social context and judgment that shape real-world decisions [65]. Qualitative data, like advisors’ meeting notes, can provide insight into “street-level” discretionary work [63, 64, 3] that is not reflected by structured data. Such data are rarely collected in tandem with quantitative measures, especially in education.
@@
Using a novel mixed-methods approach combining theoretical modeling, statistical testing, and qualitative analysis, we investigate one important form of human discretion, how MAAPS advisors use “non-algorithmic information”—information unavailable to an algorithm—to target support to students. In regular meetings, advisors actively gather information about students’ interests, struggles, and personalities, creating an information asymmetry [56] with algorithmic tools that rely on administrative data alone. By leveraging this richer context, advisors may be able to target interventions to students more effectively than would be possible using only the signals provided by any algorithm.
Our study makes two primary contributions:
Contribution 1: Theoretical Model of Expertise
@@
We develop a causal graphical framework to define and audit for human expertise in the interventional setting (§3), adding to a growing body of research studying information asymmetries in human-AI collaboration [7, 28, 31, 4, 5]. Extending prior work on auditing for human expertise in pure prediction tasks [4, 5], we define expertly targeted action in terms of three causal criteria (effectiveness, targeting, and heterogeneity) and analyze a statistical condition that is necessary but insufficient for proving expertise in interventions. We highlight a fundamental impossibility result: expertly targeted action cannot be identified from observational data alone.
Contribution 2: Quantitative and Qualitative Evidence of Expertise
@@
Using quantitative data from the MAAPS trial, we test for expertise by testing the statistical condition analyzed in our theoretical framework, whether logged advisor interventions depend on student outcomes (persistence, GPA) conditional on algorithmically available features. We identify five such potential expert interventions ($p<0.05$) (§4). Subsequent qualitative analysis of advisor comments show that advisors incorporate non-algorithmic context, such as the student’s personal circumstances, career goals, or financial troubles, into these interventions.
@@
Our findings offer evidence for the importance of advisors to the success of data-driven college advising. More broadly, they contribute to a growing body of work examining the role of human discretion in algorithm-assisted interventions.
@@ 1 Related Work
Causal inference techniques are critical for evaluating algorithm-assisted decisions and their downstream effects. Notably, Imai et al. [33] conducted the first randomized field experiment to evaluate how an algorithmic decision aid in pre-trial risk assessment affects human judgment and decision quality in the U.S. criminal justice system. Follow-up work by Ben-Michael et al. [10] focused on whether algorithm-assisted human decisions improved decision-making “accuracy” over counterfactual baselines. Our work shifts the focus from decision quality per se to the mechanisms of human discretion: we investigate whether, how, and when human decision-makers contribute additional, outcome-relevant expertise. This is especially relevant in the educational setting where advisors choose among multiple possible interventions that may improve student outcomes. Raji and Liu [55] discuss methodological challenges of evaluating algorithmic-assisted decisions arising from the cognitive responses of human decision-makers, emphasizing the need for experimental designs that disentangle the potential cognitive biases of a human judge.
In the educational domain, while the importance of understanding downstream effects of algorithmic student risk assessment has long been recognized [8], there has been little randomized evaluation of algorithmic decision making in education [40]. Perdomo et al. [52] use a regression discontinuity design to estimate the causal effect of early warning system risk scores on student success in Wisconsin public schools, finding small effects but have limited insight into underlying mechanisms due to missing data on subsequent interventions. Liu et al. [44] introduce a causal framework showing that accurate outcome prediction is insufficient for selecting effective interventions to improve individual outcomes in education.
@@
While predictive models are often used with the aim of closing equity gaps, their use can pose a risk of exacerbating racial biases in education. Notably, an investigation by The Markup found that many algorithmic predictors used by many universities incorporate race variables as predictors of student success, with effects like exacerbating historical biases, stigmatizing students, and raising risks like steering minority students into “easier” majors [20]. A host of studies demonstrate that predictive algorithms for student success sometimes make differentially accurate predictions for disadvantaged groups [51]. Additionally, scholars have analyzed other ways that algorithmic interventions in education could exacerbate inequalities beyond disparities in predictive accuracy. For instance, Baker and Hawn [6] review how bias can arise at many stages of the machine learning pipeline, from data collection to deployment. Holstein and Doroudi [30] conceptualize different ways that educational AI could exacerbate inequity by applying different “lenses” to analyze the socio-technical systems in which they are embedded.
Discretion can amplify or reduce bias. Green and Chen [26] examine how algorithmic risk scores influence human decision making in a lab study, finding that human evaluators exhibit selective adherence to algorithmic recommendations, often amplifying existing biases. Li [42] further investigates how expertise in peer review influences evaluators’ reliance on algorithmic recommendations, showing that domain experts may either mitigate or exacerbate algorithmic biases depending on their pre-existing heuristics and decision-making frameworks. Likewise, Hoffman et al. [29] analyze how hiring managers integrate algorithmic predictions into their decisions, demonstrating that discretion can either counteract or reinforce algorithmic biases. Gillis et al. [25] show that whether the algorithm amplifies or reduces the bias of the human decision-maker, with respect to a protected class, depends on whether the algorithm uses the protected class as a feature.
@@
In addition to Alur et al. [4], a growing body of work studies how humans combine their own judgment with the algorithmic predictions and what scenarios lead to humans relying more on algorithms. For example, Balakrishnan et al. [7] show that people often combine their judgments with algorithmic recommendations, increasing prediction errors by over-relying on algorithms when their own insights are more valuable and vice versa. Snyder et al. [67] find that humans rely on algorithms when the algorithms perform better than humans (on average) and when the humans are required to make decisions quickly.
@@ 2.1.2 Data
For all 1037 students (520 in treatment and 517 in control) who participated in the MAAPS study, the dataset includes their baseline characteristics at the time of enrollment to GSU (such as their demographics, high school GPA and their scores on college entry tests), as well as timeseries data on their educational outcomes updated each term (such as their GPA and credit accumulation), including their graduation outcomes. This data allows us to understand how discretionary decision making connects to students’ educational outcomes.
@@ 2.2 Our Mixed-Methods Approach
Our mixed-methods analytical framework is designed to investigate the role of advisor discretion and expertise within the treatment group. Our motivation for integrating quantitative and qualitative methods is two-fold: (1) to develop a comprehensive understanding of advisor expertise in interventional decision making, where qualitative insights help interpret and contextualize quantitative findings; and (2) to produce novel qualitative insights that can directly inform professional development for advisors, reveal gaps, and support policy recommendations.
@@
As researchers, our identities, disciplinary training, and institutional affiliations shape how we approached the research questions, interpreted the data, and interacted with the broader context of advising and algorithmic decision making. The core research team includes a graduate student and two tenure-track faculty at R1 institutions (Authors 1, 4, and 5), with training in computer science, operations research, and the societal impacts of algorithmic decision making in educational applications. The research process was shaped by a collaboration between the core research team and staff at the National Institute for Student Success (NISS) at GSU. From November 2023 to November 2024, Authors 4 and 5 held biweekly meetings with staff members at NISS, including Authors 2 and 3, to understand the MAAPS intervention and its implementation context, as well as to determine the structure and limitations of the advising data. These meetings grounded our quantitative and qualitative analyses in the realities of the program.
Our access to the MAAPS data is limited to de-identified records shared under a data use agreement with NISS, and we have not interacted directly with students or advisors. The Princeton University and Columbia University Institutional Review Boards determined the research was exempt from further review, as it does not qualify as human subjects research. Our interpretation of advising practices is shaped by our outsider perspective and by the structure and limitations of the available data. In the qualitative analysis, we sought to preserve the nuance of advisors’ written reflections while acknowledging the constraints of interpreting them at a distance. Finally, because we conduct a secondary analysis of the MAAPS experiment, the original design and data availability have also shaped our research questions and methods.
@@ 3 Theoretical Framework: Expertly Targeted Interventions
A student meets with their advisor, explaining that a medical condition prevents them from walking long distances. In response, the advisor works with the student to design a schedule with nearby classrooms that also meets their academic interests. The student is able to achieve a high semester GPA.